OpenStudy (anonymous):

h. find y’ and y” for y = 9e^((〖3x〗^2+4x-5))

OpenStudy (anonymous):

9e\[y = 9e ^{3x ^{2}}+\]

OpenStudy (anonymous):

\[y = 9e ^{(3x ^{2} + 4x - 5}\]

myininaya (myininaya):

y=9e^(g(x)) y'=9g'(x)e^(g(x)) y''=9[g''(x)e^(g(x))+g'(x)*g'(x)*e^(g(x))]

OpenStudy (anonymous):

was going to ask if you can set up few more for me

myininaya (myininaya):

ok

OpenStudy (anonymous):

i. y = ln abs value(8x^3 + 3^8x)

myininaya (myininaya):

y=ln|g(x)| y'=(g'(x))/(g(x))

myininaya (myininaya):

and if you have y=a^x lny=lna^x lny=xlna y'/y=lna y'=ylna y'=(a^x)lna so if you have y=a^x, then y'=(a^x)(lna)

OpenStudy (anonymous):

y = logbase7((5x^3 -2x^2 + 1)/ (3x^4-5))

myininaya (myininaya):

where a is a constant bigger than 0

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

do you want me to start a new thread so you can get medals.

myininaya (myininaya):

\[y=\log_7(g(x))=\frac{\ln(g(x))}{\ln7}\] it doesn't seem like equation editor is working that sucks y=logbase7(g(x))=[ln(g(x))]/[ln(7)] so y'=1/(ln7)* g'(x)/g(x)

myininaya (myininaya):

lol

myininaya (myininaya):

go ahead and make new post just in case i have to leave k? i hope you can understand the formulas i'm giving you

OpenStudy (anonymous):

trying to really hard ha

OpenStudy (anonymous):

not a math guru... =)

myininaya (myininaya):

you can do it! :)